OpenStudy (anonymous):

Ahhhhh Blue Haired Cupcake Thrower is back! I'M ALMOST DONE WITH SCHOOL! 2x+4y=14 3x+12y=51 A. (-1, 12) B. (-3, 5) C. (-1, 3) D. (-3, 20)

OpenStudy (anonymous):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

I am back too. we use the elimination rule again but at first we multiply the first equation by -3 -6x -12y = -42 3x + 12y = 51 sum them up -3x = 9 x = -3 3 * -3 + 12y = 51 -9 + 12y = 51 12 y = 60 y = 5 Check: -3 * -6 - 12 * 5 = -42 18 - 60 = -42 -42 = -42 Variant B(-3 5)

OpenStudy (anonymous):

or substitution.. 2x + 4y = 14 --- divide by 2 x + 2y = 7 x = -2y + 7 now sub -2y + 7 in for x in the other equation 3x + 12y = 51 3(-2y + 7) + 12y = 51 -6y + 21 + 12y = 51 -6y + 12y = 51 - 21 6y = 30 y = 5 now sub 5 in for y in either equation to find x 2x + 4y = 14 2x + 4(5) = 14 2x + 20 = 14 2x = 14 - 20 2x = - 6 x = -3 solution is : x = -3 and y = 5

OpenStudy (anonymous):

either way you should arrive at the same answer....whichever it is easier for you to do

OpenStudy (anonymous):

true. however in first order equations elimination is the simplest choice.